# Propagation of Interplanetary Shocks in the Heliosphere [^1] July 17, 2024

### v1

@@ 4 Analysis methods
Usually coplanarity method is applied to find the shock normal. The coplanarity method is based on the magnetic coplanarity theorem, which states that the magnetic field vectors on both sides of the shock and the shock normal lie in the same plane. Similarly, the velocity on both sides or, in other words, the velocity jump through the shock also lie in the same plane. The method of magnetic coplanarity is straightforward to implement and, it only magnetic field data is needed for applying the method (Paschmann & Daly, 1998; Schwartz, 1998).
We select two intervals upstream and downstream of the shock, calculate average upstream and downstream magnetic field vectors using the values of the interval, and determine the shock normal vector using the coplanarity method. However, in this case, the magnetic field varies strongly, therefore the calculated shock normal vector looks very sensitive for the selected intervals if one use only the coplanarity method. Hence, the calculated shock normals for each spacecraft pointed in different directions. The result cannot be explained physically, therefore applying the coplanarity method only is pointless in this case. Hence, we need the support of another method in this case.
Therefore, we use Minimum Variance Analysis (MVA) (Sonnerup & Scheible, 1998; Schwartz, 1998) and magnetic co-planarity (CP) methods to determine the IP shock normals. The MVA technique is based on the assumption that variations in the magnetic field would be observed when a single spacecraft passes through a 1D (in reality, it is a 2D or 3D transition layer) current layer or wavefront. The divergence of the magnetic field constraint $\left(\nabla\cdot\mathbf{B}=0\right)$, therefore the normal component of the magnetic field must remain constant. If such a normal direction can be found, then the variations in the magnetic field are zero or at the least have a minimum variance (Sonnerup & Cahill, 1967; Sonnerup & Scheible, 1998; Wang et al., 2024). MVA can only work if there are less fluctuations in the direction normal to the shock than in the direction perpendicular to magnetic fields upstream and downstream from the shock.
@@
    The ratio between the intermediate eigenvalue and the smallest eigenvalue should be greater than 2.
@@
We apply both methods for every single spacecraft. The shock normal vectors calculated independently point to a similar direction. This result is accaptable and could be explained physically. We suppose that the same shock passes all spacecraft at different times and locations. The shock is supposed to be a bit curved and perhaps ripped. Therefore, applying both methods leads to a valuable and acceptable result.
@@ 5 Observations
With the visual inspection of the magnetic field and plasma data, we identified jumps in magnetic field ($\mathbf{B}$), speed ($\mathbf{V}$), density (N), and temperature (T). Using these quantities, we determined the times when shock events occurred for each spacecraft data set. The common origin of a shock event detected by different spacecraft can be concluded from the similar time profiles of the magnetic field and plasma data recorded in each spacecraft position near to the shock. From the positions of the spacecraft and the times of shock arrivals to the different positions, the shock geometry and the propagation parameters (velocity, direction) can be deduced. The following case studies of two events are not orderly defined for one another. First, we analyzed the shock event on May 7, 2007, and then the event on April 23, 2007.
@@ 5.1 Event May 7, 2007
On May 7, 2007, the Wind spacecraft first detected the shock at 07:02:30 (UTC). After that, the STEREO--A spacecraft detected the shock at 08:11:30 (UTC), and then the STEREO--B spacecraft detected the shock at 09:42:00 (UTC). The four Cluster satellites detectd the shock as well. Cluster SC1 detected the shock at 08:27:55 (UTC), and Cluster SC3 detected the shock at 08:28:00 (UTC). The magnetic field and the plasma parameters are shown in Figure 1 for the Wind (a), for the STEREO--A(b), for the STEREO--B (c), in Figure 2 for the Cluster SC1 (a), and SC3 (b). There is no indication in the shock database that Cluster SC2 and SC4 detected a shock in the analysed period. However, since the four Cluster spacecraft are located close to each other, they all had to encounter the given shock. Without plasma data on Cluster SC2 and SC4 satellites, a shock cannot be confirmed based solely on the magnetic field data. However, using the empirical formula described in Section 2.4 the electron density could be estimated using EFW spacecraft potential data. Therefore, comparing the magnetic field and the electron plasma density profile, the shock is determined. The Cluster SC2 detected the shock at 08:28:10 (UTC), according to the magnetic field and density plots (Figure 2c). The Cluster SC4 detected the shock at the same time as SC2 (Figure 2d). Similarly, The electron density parameter is obtained by the empirical formula (Eq. 1).
@@
Using the results, the 2D sketches of the IP shocks that were detected by the spacecraft are shown in Figure 5. In these 2D sketches, the shock propagation and normal vector orientation are shown in a temporal development manner. Since four Cluster satellites are relatively close to one another, their averaged position as well as normal vectors are shown in the general 2D sketches (also in 3D sketches, see later). For explicitly showing normal vector directions and positions of all Clusters satellites, it is suitable to change their coordinates and normal vectors into a GSE coordinate system with the positions in the Earth radii ($R_{E}$) unit, see Figure 6.
The 3D sketches of the shock normals are shown in Figure 7a, and the 3D sketch, the STEREO--A, and B, and averaged Clusters positions are time-shifted to the position of Wind to see the overall shape of this IP shock. In order to assess the shape of the shock, we shifted the shock normals hence, as they had been detected contemporary with the shock observation time in Wind. For this, we used the average solar wind speeds and the time-differences between the observations in Wind and in the three other missions. The shifted shock normals are shown in Figure 7b. It turns out that the starting points of all of the normal vector lay approximately in a plane surface, meaning that the underlying shock exhibits a planar structure in the spatial scale of the spacecraft distances. The plane fitted to the time-shifted normals is shown in Figure 7c.
@@ 5.2 Event April 23, 2007
In this specific event, the STEREO--A spacecraft detected a shock first at 06:53:35 (UTC), then the ACE spacecraft detected the shock at 08:57:00 (UTC), and the Wind spacecraft detected the shock at 09:12:00 (UTC). The magnetic field and the plasma data are shown in Figure 8 for the STEREO--A (a), ACE (b), and Wind (c).
@@
We use all magnetic field observations of the spacecraft about the event and apply the MVA and the CP methods to them. In this event, the order between upstream and downstream is swapped because it is a fast-reverse (FR) shock event, which means the shock propagates toward the Sun against the flow however is carried outward from the Sun by the bulk solar wind flow Kilpua et al. (2015). The upstream and downstream time intervals that give the best agreement between the results of the MVA and CP shock normal determination for all spacecraft are shown in Table 4 Figure 9 shows the corresponding magnetic field observations for STEREO--A (a), ACE (b), Wind (c) and STEREO--B (d) missions. Table 4 also shows the ratio between the intermediate eigenvalue $\lambda_{2}$ and the smallest eigenvalue $\lambda_{3}$ as well as the angle between the MVA and CP normal vector determinations of the shock in the upstream and downstream time intervals for each spacecraft.
@@
Similarly to the Event May 7, the upstream and downstream time intervals are defined correctly as well, considering the angle difference between the two methods is minimal and the eigenvalue criteria are fulfilled for each spacecraft data. Also, using the determined time intervals, in Table 5 we summarize, for each mission, the ratio between the upstream and downstream total magnetic fields, densities and temperatures, the differences between the upstream and downstream solar wind bulk speeds. From the obtained upstream/downstream ratios and bulk speed differences we conclude that the criteria for detecting a sock event in the considered period is fulfilled (Lumme et al., 2017) for each spacecraft. Even for STEREO--B, for which shock was not indicated in the shock catalogue (see above), the numbers unambiguously prove the existence of a shock event during the studied interval. However, the almost quasi-parallel shock detected by the STEREO--B appears to become quasi-parallel a few hours after the shock detection times of STEREO--A, ACE, and the Wind based on the shock $\theta_{Bn}$ angle, see Table 5, where the calculated results of additional parameters are shown.
@@
The sketches of the IP shocks geometries that were detected by the spacecraft are shown in Figure 10. In these 2D sketches, the shock propagation and normal vector orientation are shown in a temporal development manner.
@@
The 3D sketch of the IP shock geometries are shown in Figure 11. The 3D sketch, the STEREO--A, and B positions are time-shifted to the position of STEREO--A to see the overall shape of this IP shock. In order to assess the shape of the shock, we shifted the shock normals hence, as they had been detected contemporary with the shock observation time in STEREO--A. For this, we used the average solar wind speeds and the time-differences between the observations in STEREO--A and in the two other spacecraft. The shifted shock normals are shown in the top right panel of Figure 11. The top view of the time-shifted normals is shown in the bottom part of Figure 11.
@@ 6 Discussion
The IP shock parameters did not change significantly most separated couple of probes, from STEREO--A to B in neither case (Table 2, 35, 6). There is no sign of wave dispersion on the scale of 40 million km, that is the distance of the STEREO spacecraft.
@@ 6.1 Event May 7, 2007
The IP shock fitted plane is tilted $56.42^{\circ}$ from the Sun-Earth line according to Figure 7c. The tilt appears to be almost the same as expected for the Parker spiral impacting the Earth from the dawn side. Hence, it is the reason why the Wind detected the shock first even though STEREO--A’s position is relatively closer to the Sun’s direction. This period, 2007, was during the solar minimum phase, and stream interaction region (SIR) or co-rotating stream interactions (CIR) were dominant (Opitz et al., 2014), and this further proves the result, see Figure 12. Furthermore, to see a correlation between fast-forward shock and geomagnetic activity, we investigated the $K_{p}$-index (Bartels, 1949; Rostoker, 1972; Matzka et al., 2021, https://www.swpc.noaa.gov/noaa-scales-explanation) on May 07, 2007, as shown in Figure 13. It appears that this forward shock-leading event disturbed the magnetosphere, causing a G1-minor geomagnetic storm with the $K_{p}$- index peaking around 15:00 (UTC). The geomagnetic substorm happened about 6 hours after the detection of the shock by the STEREO--B spacecraft.
@@ 6.2 Event April 23, 2007
In Figure 11 it seems that the shape of the shock is not uniform and is twisted from the STEREO--A spacecraft to the other three spacecraft in the Z-axis along the transverse direction (Y-axis) to the Sun-Earth line, see Figure 11b. As seen from Figure 11c, the STEREO--A alone is on the dusk (left) side of the Sun-Earth line while the Wind, the ACE, and the STEREO--B are on the dawn (right). In this XY-plane view (Figure 11), the shock normal vectors appear to be changing or slightly rotating their direction from the STEREO--A to the Wind, the ACE, and the STEREO--B about the Z-axis, and the shock is changing from the quasi-perpendicular to quasi-parallel based on the shock $\theta_{Bn}$ angle. This IP shock is a fast reverse shock, meaning it travels toward the Sun in a frame fixed to the solar wind even though the shock propagates away from the Sun with the solar wind. The shock detection time difference between the STEREO--A and the Wind/ACE is about two hours while between STEREO--A and B is almost six hours, yet the shock orientation is significantly changed from the STEREO--A to the other three spacecraft indicating the change is spatial, not temporal. So, due to this nature, the IP shock could be a local ripple. The ripples on the shock surface are known be caused by ICME (interplanetary coronal mass ejections) shock drivers as they do not propagate into homogeneous interplanetary medium (Acuña et al., 2008). However, for this particular event the source is suggested to be a stream interaction region (SIR) according to the catalogue given in https://stereo-ssc.nascom.nasa.gov/pub/ins_data/impact/level3/STEREO_Level3_Shock.pdf.
@@
We investigated the $K_{p}$-index of this event as seen in Figure 14. Similar to the fast-forward shock event of May 07, 2007, a G1-minor geomagnetic storm occurred on this day (see space weather scales of National Oceanic and Atmospheric Administration (NOAA) https://www.swpc.noaa.gov/noaa-scales-explanation), but the beginning of the geomagnetic storm happened three hours before the shock detection time of the STEREO--A while the ending of the storm was partially at the same time as the STEREO--A. As stated before, SIR/CIRs form when the fast-moving solar wind catches the slow-moving solar wind Richardson (2018). This sometimes forms a pair of shocks, one leading as a fast-forward shock while a rarefied shock trails the solar wind as a fast-reverse shock but oftentimes as just sole fast-forward or fast reverse shocks (Jian et al., 2019). Nevertheless, since it always involves the fast-moving solar wind, it makes total sense why the G1-minor geomagnetic storm with a $K_{p}$-5 index happened before the detection of the shock because it seems the fast-moving solar wind caused the minor storm before the fast-reverse shock finally arrived at the spacecraft positions.
@@
The normal orientation inconsistency can also be explained by the argument that this shock is a fast-reverse shock because SIRs have characteristic tilts such that forward waves direct towards the solar equatorial plane while the reverse waves tend to move in the direction of the solar poles (Kilpua et al., 2015). Therefore, this tilting nature may explain the XZ component tilted orientation between the STEREO--A and the other three spacecraft – the Wind, the ACE, and the STEREO--B.

### v2

@@ 4 Analysis methods
Usually coplanarity method is applied to find the shock normal. The coplanarity method is based on the magnetic coplanarity theorem, which states that the magnetic field vectors on both sides of the shock and the shock normal lie in the same plane. Similarly, the velocity on both sides or, in other words, the velocity jump through the shock also lie in the same plane. The method of magnetic coplanarity is straightforward to implement, it requires only magnetic field data for applying the method (Paschmann & Daly, 1998; Schwartz, 1998).
We select two intervals upstream and downstream of the shock, calculate average upstream and downstream magnetic field vectors using the values of the interval, and determine the shock normal vector using the coplanarity method. However, in this case, the magnetic field varies strongly, therefore the calculated shock normal vector looks very sensitive for the selected intervals if one uses only the coplanarity method. Hence, the calculated shock normals for each spacecraft pointed in different directions. The result cannot be explained physically, therefore applying the coplanarity method only is pointless in this case. Hence, we need the support of another method in this case.
Therefore, we use Minimum Variance Analysis (MVA) (Sonnerup & Scheible, 1998; Schwartz, 1998) and magnetic co-planarity (CP) methods to determine the IP shock normals. The MVA technique is based on the assumption that variations in the magnetic field would be observed when a single spacecraft passes through a 1D (in reality, it is a 2D or 3D transition layer) current layer or wavefront. The divergence of the magnetic field is constrained $\left(\nabla\cdot\mathbf{B}=0\right)$, therefore the normal component of the magnetic field must remain constant. If such a normal direction can be found, then the variations in the magnetic field are zero or at least have a minimum variance (Sonnerup & Cahill, 1967; Sonnerup & Scheible, 1998; Wang et al., 2024). MVA can only work if there are less fluctuations in the direction normal to the shock than in the direction perpendicular to magnetic fields upstream and downstream from the shock.
@@
    The ratio between the intermediate eigenvalue and the smallest eigenvalue should be greater than 2 in the MVA method.
@@
We apply both methods for every single spacecraft. The shock normal vectors calculated independently point to a similar direction. This result is acceptable and could be explained physically. We suppose that the same shock passes all spacecraft at different times and locations. The shock is supposed to be a bit curved and perhaps rippled. Therefore, applying both methods leads to a valuable and acceptable result.
@@ 5 Observations
With the visual inspection of the magnetic field and plasma data, we identified jumps in magnetic field ($\mathbf{B}$), speed ($\mathbf{V}$), density (N), and temperature (T). Using these quantities, we determined the times when shock events occurred for each spacecraft data set. The common origin of a shock event detected by different spacecraft can be concluded from the similar time profiles of the magnetic field and plasma data recorded in each spacecraft position near the shock. From the positions of the spacecraft and the times of shock arrivals to the different positions, the shock geometry and the propagation parameters (velocity, direction) can be deduced. The following case studies of two events are discussed in reversed time order. First, we analyzed the shock event on May 7, 2007, and then the event on April 23, 2007.
@@ 5.1 Event May 7, 2007
On May 7, 2007, the Wind spacecraft detected the shock first at 07:02:30 (UTC). After that, the STEREO--A spacecraft detected the shock at 08:11:30 (UTC), and then the STEREO--B spacecraft detected the shock at 09:42:00 (UTC). The four Cluster satellites detected the shock as well. Cluster SC1 detected the shock at 08:27:55 (UTC), and Cluster SC3 detected the shock at 08:28:00 (UTC). The magnetic field and the plasma parameters are shown in Figure 1 for the Wind (a), for the STEREO--A(b), for the STEREO--B (c), in Figure 2 for the Cluster SC1 (a), and SC3 (b). There is no indication in the shock database that Cluster SC2 and SC4 detected a shock in the analysed period. However, since the four Cluster spacecraft are located close to each other, they all had to encounter the given shock. Without plasma data on Cluster SC2 and SC4 satellites, a shock cannot be confirmed based solely on the magnetic field data. However, using the empirical formula described in Section 2.4 the electron density could be estimated using EFW spacecraft potential data. Therefore, comparing the magnetic field and the electron plasma density profile, the shock is determined. The Cluster SC2 detected the shock at 08:28:10 (UTC), according to the magnetic field and density plots (Figure 2c). The Cluster SC4 detected the shock at the same time as SC2 (Figure 2d). Similarly, The electron density parameter is obtained by the empirical formula (Eq. 1).
@@
Using the results, the 2D sketches of the IP shocks that were detected by the spacecraft are shown in Figure 5. In these 2D sketches, the shock propagation and normal vector orientation are shown in a temporal development manner. Since four Cluster satellites are relatively close to one another, their averaged position as well as normal vectors are shown in the general 2D sketches (also in 3D sketches, see later). For explicitly showing normal vector directions and positions of all Cluster satellites, it is suitable to change their coordinates and normal vectors into a GSE coordinate system with the positions in the Earth radii ($R_{E}$) unit, see Figure 6.
The 3D sketches of the shock normals are shown in Figure 7. In order to assess the shape of the shock, we shifted the shock normals hence, as they had been detected contemporary with the shock observation time in Wind. For this, we used the average solar wind speeds and the time differences between the observations in Wind and the three other missions. The shifted shock normals are shown in Figure 7b. It turns out that the starting points of all of the normal vectors lay approximately in a plane surface, meaning that the underlying shock exhibits a planar structure in the spatial scale of the spacecraft distances. The plane fitted to the time-shifted normals is shown in Figure 7c.
@@ 5.2 Event April 23, 2007
In this specific event, the STEREO--A spacecraft detected a shock first at 06:53:35 (UTC), then the ACE spacecraft detected the shock at 08:57:00 (UTC), and the Wind spacecraft detected the shock at 09:12:00 (UTC). The magnetic field and the plasma data are shown in Figure 8 for the STEREO--A (a), ACE (b), Wind (c), and STEREO--B (d).
@@
We use all magnetic field observations of the spacecraft about the event and apply the MVA and the CP methods to them. In this event, the order between upstream and downstream is swapped because it is a fast-reverse (FR) shock event, which means the shock propagates toward the Sun against the flow but is carried outward from the Sun by the bulk solar wind flow Kilpua et al. (2015). The upstream and downstream time intervals that give the best agreement between the results of the MVA and CP shock normal determination for all spacecraft are shown in Table 4. Figure 9 shows the corresponding magnetic field observations for STEREO--A (a), ACE (b), Wind (c) and STEREO--B (d) missions. Table 4 also shows the ratio between the intermediate eigenvalue $\lambda_{2}$ and the smallest eigenvalue $\lambda_{3}$ as well as the angle between the MVA and CP normal vector determinations of the shock in the upstream and downstream time intervals for each spacecraft.
@@
Similarly to the Event May 7, the upstream and downstream time intervals are defined correctly as well, considering the angle difference between the two methods is minimal and the eigenvalue criteria are fulfilled for each spacecraft data. Also, using the determined time intervals, in Table 5 we summarize, for each mission, the ratio between the upstream and downstream total magnetic fields, densities and temperatures, and the differences between the upstream and downstream solar wind bulk speeds. From the obtained upstream/downstream ratios and bulk speed differences we conclude that the criteria for detecting a shock event in the considered period is fulfilled (Lumme et al., 2017) for each spacecraft. Even for STEREO--B, for which shock was not indicated in the shock catalogue (see above), the numbers unambiguously prove the existence of a shock event during the studied interval. However, the almost quasi-parallel shock detected by the STEREO--B appears to become quasi-parallel a few hours after the shock detection times of STEREO--A, ACE, and the Wind based on the shock $\theta_{Bn}$ angle, see Table 5, where the calculated results of additional parameters are shown.
@@
The sketches of the IP shock geometries that were detected by the spacecraft are shown in Figure 10. In these 2D sketches, the shock propagation and normal vector orientation are shown in a temporal development manner.
@@
The 3D sketch of the IP shock geometries is shown in Figure 11. The 3D sketch, the STEREO--A, and B positions are time-shifted to the position of STEREO--A to see the overall shape of this IP shock. In order to assess the shape of the shock, we shifted the shock normals hence, as they had been detected contemporary with the shock observation time in STEREO--A. For this, we used the average solar wind speeds and the time differences between the observations in STEREO--A and in the two other spacecraft. The shifted shock normals are shown in the top right panel of Figure 11. The top view of the time-shifted normals is shown in the bottom part of Figure 11.
@@ 6 Discussion
The IP shock parameters did not change significantly from STEREO--A to B in either case (Table 2, 35, 6). The IP shocks did not evolve on the scale of 40 million km, that is the distance of the STEREO spacecraft.
@@ 6.1 Event May 7, 2007
The IP shock fitted plane is tilted $56.42^{\circ}$ from the Sun-Earth line according to Figure 7c. The tilt appears to be almost the same as expected for the Parker spiral impacting the Earth from the dawn side. Hence, it is the reason why the Wind detected the shock first even though STEREO--A’s position is relatively closer to the Sun’s direction. This period, 2007, was during the solar minimum phase, and stream interaction region (SIR) or corotating interaction region (CIR) were dominant (Opitz et al., 2014), and this further proves the result, see Figure 12. Furthermore, to see a correlation between fast-forward shock and geomagnetic activity, we investigated the $K_{p}$-index (Bartels, 1949; Rostoker, 1972; Matzka et al., 2021, https://www.swpc.noaa.gov/noaa-scales-explanation) on May 07, 2007, as shown in Figure 13. It appears that this forward shock-leading event disturbed the magnetosphere, causing a G1-minor geomagnetic storm with the $K_{p}$- index peaking around 15:00 (UTC). The geomagnetic substorm happened about 6 hours after the detection of the shock by the STEREO--B spacecraft.
@@ 6.2 Event April 23, 2007
In Figure 11 it seems that the shape of the shock is not uniform and is twisted from the STEREO--A spacecraft to the other three spacecraft about the Z-axis along the transverse direction (Y-axis) to the Sun-Earth line, see Figure 11b. As seen from Figure 11c, the STEREO--A alone is on the dawn (left) side of the Sun-Earth line while the Wind, the ACE, and the STEREO--B are on the dusk (right). In this XY-plane view (Figure 11), the shock normal vectors appear to be changing or slightly rotating their direction from the STEREO--A to the Wind, the ACE, and the STEREO--B about the Z-axis, and the shock is changing from the quasi-perpendicular to quasi-parallel based on the shock $\theta_{Bn}$ angle. This IP shock is a fast reverse shock, meaning it travels toward the Sun in a frame fixed to the solar wind even though the shock propagates away from the Sun with the solar wind. The shock detection time difference between the STEREO--A and the Wind/ACE is about two hours while between STEREO--A and B is almost six hours. Yet the shock orientation is significantly changed from the STEREO--A to the other three spacecraft indicating the change is spatial, not temporal. So, due to this nature, the IP shock could be a local ripple. The ripples on the shock surface are known to be caused by ICME (interplanetary coronal mass ejections) shock drivers as they do not propagate into homogeneous interplanetary medium (Acuña et al., 2008). However, for this particular event the source is suggested to be a stream interaction region (SIR) according to the catalogue given in https://stereo-ssc.nascom.nasa.gov/pub/ins_data/impact/level3/STEREO_Level3_Shock.pdf.
@@
We investigated the $K_{p}$-index of this event as seen in Figure 14. Similar to the fast-forward shock event of May 07, 2007, a G1-minor geomagnetic storm occurred on this day (see space weather scales of National Oceanic and Atmospheric Administration (NOAA) https://www.swpc.noaa.gov/noaa-scales-explanation), but the beginning of the geomagnetic storm happened three hours before the shock detection time of the STEREO--A while the ending of the storm was partially at the same time as the STEREO--A shock detection. As stated before, SIR/CIRs form when the fast-moving solar wind catches the slow-moving solar wind Richardson (2018). This sometimes forms a pair of shocks, one leading as a fast-forward shock while a rarefied shock trails the solar wind as a fast-reverse shock but oftentimes as just sole fast-forward or fast reverse shocks (Jian et al., 2019). Nevertheless, since it always involves the fast-moving solar wind, it makes total sense why the G1-minor geomagnetic storm with a $K_{p}=5$ index happened before the detection of the shock because it seems the fast-moving solar wind caused the minor storm before the fast-reverse shock finally arrived at the spacecraft positions.
@@
The normal orientation inconsistency can also be explained by the argument that this shock is a fast-reverse shock because SIRs have characteristic tilts such that forward waves direct towards the solar equatorial plane while the reverse waves tend to move in the direction of the solar poles (Kilpua et al., 2015). Therefore, this tilting nature may explain the XZ component’s tilted orientation between the STEREO--A and the other three spacecraft – the Wind, the ACE, and the STEREO--B.